### v2

@@ Abstract
While there already exist randomized subspace Newton methods that restrict the search direction to a random subspace for a convex function, we propose a randomized subspace regularized Newton method for a non-convex function and more generally we investigate thoroughly the local convergence rate of the randomized subspace Newton method. In our proposed algorithm using a modified Hessian of the function restricted to some random subspace, with high probability, the function value decreases even when the objective function is non-convex. In this paper, we show that our method has global convergence under appropriate assumptions and its convergence rate is the same as that of the full regularized Newton method. Furthermore, we can obtain a local linear convergence rate under some additional assumptions, and prove that this rate is the best we can hope, in general, when using random subspace. We furthermore prove that if the Hessian at the local optimum is rank defficient then superlienar convergence holds.
@@ 1 Introduction
$f_{x_{k}}\ :\ u\mapsto f(x_{k}+P_{k}^{\top}u).$ (2)
Other second-order subspace descent methods, as well cubically-regularized subspace Newton methods, [13], have been studied in the literature. More precisely, the method in [13] can be seen as a stochastic extension of the cubically-regularized Newton method [21] and also as a second-order enhancement of stochastic subspace descent [17]. In [16], a random subspace version of the BFGS method is proposed. The authors prove local linear convergence if the function is assumed to be self-concordant. Apart in recent Shao’s Ph.D thesis [24] and the associated papers [9, 8] which have been done parallelly to this paper, to the best of our knowledge, existing second-order subspace methods have iteration complexity analysis only for convex optimization problems.
The thesis [24] and the paper [8] propose a random subspace adaptive regularized cubic method for unconstrained non-convex optimization and show a global convergence property with sub-linear rate to a stationary point [^2] The author also proves that if the Hessian matrix has low rank and scaled Gaussian sketching matrices are used, then the Hessian at the stationary point is approximately positive semidefinite with high probability.. In this paper we propose a new subspace method based on the regularized Newton method and discuss the local convergence rate together with global iteration complexity. [^3] Just as the ordinary cubic method is superior to the Newton method in terms of iteration complexity, similar observation seems to hold between the subspace cubic method [24] and ours. Notice indeed that, to the best of our knowledge, the local convergence of such methods never seen to have been thoroughly studied [^4] Some papers, as we will see later, investigate when local linear convergence holds.; one would expect super-linear convergence for second order method and no papers discuss whether super-linear convergence holds or not for second order methods. Indeed any iterative algorithm can be easily adapted to a random subspace method as it suffices to apply it to the function restricted to the subspace: $u\mapsto f(x_{k}+P^{\top}_{k}u)$. We therefore believe that it is important to investigate thoroughly whether the properties of such full-space algorithms are preserved or not when adapted to the random subspace setting.
The topic around subspace methods in optimization has received much attention recently. Besides random subspace methods, deterministic subspace methods are also investigated. In [20], a limited-memory common-directions method is proposed for smooth optimization problems. At each iteration, a subspace is constructed using first-order information from previous iterations, and a Newton method is used to obtain an approximate minimizer within this subspace. They prove sub-linear convergence in the non-convex case and linear convergence if the function satisfies the Polyak-Lojasiewicz (PL) inequality:
$\forall x\in\mathbb{R}^{n},\quad\|\nabla f(x)\|^{2}\geq c_{0}(f(x)-f(x^{*})).$ (3)
@@
where $\eta_{k}$ is defined to ensure that search direction $d_{k}$ is a descent direction and the step size $t_{k}$ is chosen so that it satisfies Armijo’s rule. As with RSN, this algorithm is expected to be computationally efficient since we use projections onto lower-dimensional spaces. In this paper, we first show that RS-RNM has global convergence under appropriate assumptions, more precisely, we have $\left\|\nabla f(x_{k})\right\|\leq\varepsilon$ after at most $O(\varepsilon^{-2})$ iterations with some probability, which is the same as the global convergence rate shown in [26] for the full regularized Newton method. We then prove that under additional assumptions, we can obtain a linear convergence rate locally. In particular, one contribution of the paper is to propose, to the best of our knowledge, the weakest conditions until now for local linear convergence. To do so we will extensively use the fact that the subspace is chosen at random. From these conditions, we can derive a random-projection version of the PL inequality (3), which will be satisfied when the function is restricted to a random subspace. One other contribution of this paper is to prove that, in general, linear convergence is the best rate we can hope for this method. Furthermore, we also prove that if the Hessian at the local optima is rank deficient, then one can achieve super-linear convergence using a subspace dimension $s$ large enough.
@@ 2.2 Properties of random projection
Lemma 2.1 ([28, Lemma 5.3.2, Exercise 5.3.3]).
@@
Proof. By [28, Theorem 4.6.1], there exists a constant $\bar{C}$ such that
@@
For any two events $\mbox{For any two events }E_{1}\mbox{ and }E_{2}:\ \rm{Prob}(E_{1}\cap E_{2})\geq 1-\left((1-\rm{Prob}(E_{1}))+(1-\rm{Prob}(E_{2}))\right).$ (4)
@@ 3 Randomized subspace regularized Newton method
In [26], the authors develop one of the regularized Newton methods (RNM):
@@
where $\Lambda_{k}=\max(0,-\lambda_{\min}(H_{k}))$, $c_{1}^{\prime},c_{2}^{\prime},\gamma^{\prime}$ are some positive parameter values and $t_{k}$ is the step-size chosen by Armijo’s step size rule, and show that this algorithm achieves $\left\|g_{k}\right\|\leq\varepsilon$ after at most $O(\varepsilon^{-2})$ iterations and it has a super-linear rate of convergence in a neighborhood of a local optimal solution under appropriate conditions.
@@
Let $\mathcal{D}$ denotes the set of Gaussian matrices of size $s\times n$ whose entries are independently sampled from $\mathrm{N}(0,1/s)$. With a Gaussian random matrix $P_{k}$ from $\mathcal{D}$, the regularized sketched Hessian is defined by:
$M_{k}:=P_{k}H_{k}P_{k}^{\mathsf{T}}+\eta_{k}I_{s}\in\mathbb{R}^{s\times s},$ (6)
@@
$d_{k}:=-P_{k}^{\mathsf{T}}M_{k}^{-1}P_{k}g_{k}.$ (7)
The costly part of Newton-based methods, the inverse computation of a (approximate) Hessian matrix, is done in the subspace of size $s$. We note that $d_{k}$ defined by 7 is a descent direction for $f$ at $x_{k}$, i.e., $g_{k}^{\top}d_{k}<0$ if $g_{k}\neq 0$, since it turns out that $M_{k}$ is positive definite from the definition of $\Lambda_{k}$, and therefore $P_{k}^{\mathsf{T}}M_{k}^{-1}P_{k}$ is also positive definite (with high probability).
@@
$f(x_{k})-f(x_{k}+\beta^{l_{k}}d_{k})\geq-\alpha\beta^{l_{k}}g_{k}^{\mathsf{T}}d_{k}.$ (8)
@@ 4 Global convergence properties
In Section 4.1, we discuss the global convergence of the RS-RNM under Assumptions: 4. We further prove the global iteration complexity of the algorithm in Section 4.2 by assuming further assumptions.
@@
By (8), we have that for any $k\in\mathbb{N}$, $f(x_{k+1})\leq f(x_{k})$, implying all $x_{k}\in\Omega$. Since $\Omega$ is a bounded set and $f$ is continuously differentiable, there exists $U_{g}>0$ such that
$\left\|g_{k}\right\|\leq U_{g},\ \forall k\geq 0.$ (9)
@@
$\|H_{k}\|\leq L.$ (11)
Notice that by (10), $\nabla f$ is $L$-Lipschitz continuous. We also define $f^{*}=\inf_{x\in\Omega}f(x)$.
@@ 4.1 Global convergence
Suppose that $\left\|d_{k}\right\|\neq 0$. Then, $d_{k}$ defined by 7 satisfies
$\left\|d_{k}\right\|\leq\mathcal{C}\frac{n}{s}\frac{\left\|g_{k}\right\|^{1-\gamma}}{c_{2}},$
@@
Proof. By Lemma 2.2 we have $\left\|P_{k}^{\mathsf{T}}P_{k}\right\|\leq\mathcal{C}\frac{n}{s}$, holds with probability at least $1-2e^{-s}$. Then, it follows from 7 that
@@
We next show that, when $\left\|g_{k}\right\|$ is at least $\varepsilon$ away from 00, $\left\|d_{k}\right\|$ is bounded above by some constant.
@@
Suppose that Assumption 4 holds. Suppose also that there exists $\varepsilon>0$ such that $\left\|g_{k}\right\|\geq\varepsilon$. Then, with probability at least $1-2e^{-s}$, $d_{k}$ defined by 7 satisfies
$\left\|d_{k}\right\|\leq r(\varepsilon),$ (12)
@@
Proof. If $\gamma\leq 1$, it follows from Lemma 4.1 and 9 that
@@
$\left\|\nabla^{2}f(x)\right\|\leq U_{H}(\varepsilon),\ \forall x\in\Omega+B(0,r(\varepsilon)).$ (13)
@@
Suppose that Assumption 4 holds. Suppose also that there exists $\varepsilon>0$ such that $\left\|g_{k}\right\|\geq\varepsilon$. Then, with probability at least $1-2e^{-s}$, a step size $t_{k}^{\prime}>0$ such that
@@
By Lemma 2.2, we have that, with probability at least $1-2e^{-s}$, $\left\|P_{k}P_{k}^{\mathsf{T}}\right\|\leq\frac{\bar{\mathcal{C}}n}{s}$.
$f(x_{k})-f(x_{k}+t_{k}^{\prime}d_{k})+\alpha t_{k}^{\prime}g_{k}^{\mathsf{T}}d_{k}$
$=$ $(1-\alpha)t_{k}^{\prime}g_{k}^{\mathsf{T}}P_{k}^{\mathsf{T}}M_{k}^{-1}P_{k}g_{k}-\frac{1}{2}{t_{k}^{\prime}}^{2}g_{k}^{\mathsf{T}}P_{k}^{\mathsf{T}}M_{k}^{-1}P_{k}\nabla^{2}f(x_{k}+\tau_{k}^{\prime}t_{k}^{\prime}d_{k})P_{k}^{\mathsf{T}}M_{k}^{-1}P_{k}g_{k}$ (14)
(by 7)
[…]
(by 13)
In addition, we have $\left\|H_{k}\right\|\leq U_{H}(\varepsilon)$ from 13, which gives us $\left\|P_{k}H_{k}P_{k}^{\mathsf{T}}\right\|\leq\frac{\bar{\mathcal{C}}n}{s}U_{H}(\varepsilon)$. For these reasons, we obtain evaluation of the values of $\lambda_{\min}(M_{k}^{-1})$ and $\lambda_{\max}(M_{k}^{-1}P_{k}P_{k}^{\mathsf{T}}M_{k}^{-1})$:
@@
$\geq\frac{1}{\frac{\bar{\mathcal{C}}n}{s}U_{H}(\varepsilon)+c_{1}\frac{\bar{\mathcal{C}}n}{s}U_{H}(\varepsilon)+c_{2}\left\|g_{k}\right\|^{\gamma}},$ (15)
@@
(by 9 and $\left\|g_{k}\right\|\geq\varepsilon$)
@@
Suppose that Assumption 4 holds. Suppose also that there exists $\varepsilon>0$ such that $\left\|g_{k}\right\|\geq\varepsilon$. Then, with probability at least, $1-2e^{-s}$ the step size $t_{k}$ chosen in Line 6 of RS-RNM satisfies
$t_{k}\geq t_{\min}(\varepsilon),$ (16)
@@
we know that $t_{k}=1$ satisfies Armijo’s rule 8 from 4.5. If not, there exists $l_{k}\in\{0,1,2,\ldots\}$ such that
@@
and by Lemma 4.5, we have that the step size $\beta^{l_{k}+1}$ satisfies Armijo’s rule 8. Then, from the definition of $\beta^{l_{k}}$ in Line 6 of RS-RNM, we have
@@
Using Corollary 4.7, we can show the global convergence of RS-RNM under Assumptions 4.
@@
Suppose that Assumption 4 holds. Let
[…]
Then, with probability at least $1-2m\left(\exp(-\frac{{\mathcal{C}}_{0}}{4}s)+\exp(-s)\right)$, there exists $k\in\{0,1,\dots,m-1\}$ such that $\left\|g_{k}\right\|<\varepsilon$.
@@
Proof. We first notice that, by Lemma 2.1, applied with $\varepsilon=1/2$, and Lemma 2.2, we have, using (4), that $\left\|P_{k}g_{k}\right\|^{2}\geq\frac{1}{2}\left\|g_{k}\right\|^{2}$ and $\|P_{k}P_{k}^{\top}\|\leq\bar{\mathcal{C}}\frac{n}{s}$ holds for all $k\in\{0,1,\ldots,m-1\}$ with the given probability.
Suppose, for the sake of contradiction, that $\left\|g_{k}\right\|\geq\varepsilon$ for all $k\in\{0,1,\dots,m-1\}$. From Armijo’s rule 8, we can estimate how much the function value decreases in one iteration:
@@
Adding up all these inequalities from $k=0$ to $k=m-1$, we get
$f(x_{0})-f(x_{m})\geq mp(\varepsilon)\varepsilon^{2},$
and thus,
@@
$\geq$ $mp(\varepsilon)\varepsilon^{2}$
$>$ $\frac{f(x_{0})-f^{*}}{p(\varepsilon)\varepsilon^{2}}\cdot p(\varepsilon)\varepsilon^{2}$
$=$ $f(x_{0})-f^{*},$
which is a contradiction.

### v3

@@ Abstract
While there already exist randomized subspace Newton methods that restrict the search direction to a random subspace for a convex function, we propose a randomized subspace regularized Newton method for a non-convex function and more generally we investigate thoroughly, for the first time, the local convergence rate of the randomized subspace Newton method. In our proposed algorithm, we use a modified Hessian of the function restricted to some random subspace so that, with high probability, the function value decreases at each iteration, even when the objective function is non-convex. We show that our method has global convergence under appropriate assumptions and its convergence rate is the same as that of the full regularized Newton method. Furthermore, we obtain a local linear convergence rate under some additional assumptions, and prove that this rate is the best we can hope, in general, when using a random subspace. We furthermore prove that if the Hessian, at the local optimum, is rank deficient then super-linear convergence holds.
@@ 1 Introduction
(2) $f_{x_{k}}\ :\ u\mapsto f(x_{k}+P_{k}^{\top}u).$
Other second-order subspace descent methods, such as cubically-regularized subspace Newton methods, [22], have been studied in the literature. More precisely, the method in [22] can be seen as a stochastic extension of the cubically-regularized Newton method [32] and also as a second-order enhancement of stochastic subspace descent [28]. In [27], a random subspace version of the BFGS method is proposed. The authors prove local linear convergence, if the function is assumed to be self-concordant. Apart in recent Shao’s Ph.D thesis [37] and the associated papers [12, 11] which have been done parallelly to this paper, to the best of our knowledge, existing second-order subspace methods have iteration complexity analysis only for convex optimization problems.
The thesis [37] and the paper [11] propose a random subspace adaptive regularized cubic method for unconstrained non-convex optimization and show a global convergence property with sub-linear rate to a stationary point [^2] The author also proves that if the Hessian matrix has low rank and scaled Gaussian sketching matrices are used, then the Hessian at the stationary point is approximately positive semidefinite with high probability.. In this paper we propose a new subspace method based on the regularized Newton method and discuss the local convergence rate together with global iteration complexity. [^3] Just as the ordinary cubic method is superior to the Newton method in terms of iteration complexity, similar observation seems to hold between the subspace cubic method [37] and ours. Notice indeed that, to the best of our knowledge, the local convergence of such methods never seems to have been thoroughly studied [^4] Some papers, as we will see later, investigate when local linear convergence holds.; one would expect super-linear convergence for second order methods and no papers discuss whether super-linear convergence holds or not for second order methods. Indeed any iterative algorithm can be easily adapted to a random subspace method as it suffices to apply it to the function restricted to the subspace: $u\mapsto f(x_{k}+P^{\top}_{k}u)$. We therefore believe that it is important to investigate thoroughly whether the properties of such full-space algorithms are preserved or not when adapted to the random subspace setting.
@@
where $\eta_{k}$ is defined to ensure that search direction $d_{k}$ is a descent direction and the step size $t_{k}$ is chosen so that it satisfies Armijo’s rule. As with RSN, this algorithm is expected to be computationally efficient since we use projections onto lower-dimensional spaces. In this paper, we first show that RS-RNM has global convergence under appropriate assumptions, more precisely, we have $\left\|\nabla f(x_{k})\right\|\leq\varepsilon$ after at most $O(\varepsilon^{-2})$ iterations with some probability, which is the same as the global convergence rate shown in [39] for the full regularized Newton method. We then prove that under additional assumptions, we can obtain a linear convergence rate locally. In particular, one contribution of the paper is to propose, to the best of our knowledge, the weakest conditions until now for local linear convergence. To do so we will extensively use the fact that the subspace is chosen at random. From these conditions, we can derive a random-projection version of the Polyak-Lojasiewicz (PL) inequality (3),
(3) $\forall x\in\mathbb{R}^{n},\quad\|\nabla f(x)\|^{2}\geq c_{0}(f(x)-f(x^{*})),$
which will be satisfied when the function is restricted to a random subspace. One other contribution of this paper is to prove that, in general, linear convergence is the best rate we can hope for this method. Furthermore, we also prove that if the Hessian at the local optima is rank deficient, then one can achieve super-linear convergence using a subspace dimension $s$ large enough.
Our randomized subspace method for nonconvex optimization problems is based on the regularized Newton method in [39, 40]. While various other regularized Newton methods have been proposed in recent years, most of them are for convex problems or non-smooth optimization problems. For example, [31] presents a globally convergent proximal Newton-type method for non-smooth convex optimization and [8] develops coderivative-based Newton methods combined with Wolfe line-search for non-smooth problems. Recently [45] proposes a generalized regularization method that includes quadratic, cubic, and elastic net regularizations. Also [14] proposes, in the convex case, a variant of the Newton method with quadratic regularization and proves better global rate. Recent papers, [19, 47, 48], propose regularization methods for the non-convex case. However, although these methods obtained better iterations complexity, the subroutines involved to compute are quite complex and not as simple as in [39, 40]. By applying similar random subspace techniques to these methods, we may be able to develop random subspace variants with state-of-the-art theoretical guarantees, but that is a topic for future work.
@@ 2.2 Properties of random projection
Lemma 2.1 ([41, Lemma 5.3.2, Exercise 5.3.3]).
@@
Proof. By [41, Theorem 4.6.1], there exists a constant $\bar{C}$ such that
@@
(4) $\mbox{For any two events }E_{1}\mbox{ and }E_{2}:\ \rm{Prob}(E_{1}\cap E_{2})\geq 1-\left((1-\rm{Prob}(E_{1}))+(1-\rm{Prob}(E_{2}))\right).$
@@ 3 Randomized subspace regularized Newton method
The paper [39] develops a regularized Newton methods (RNM) that constructs a sequence of iterates with the following update rule:
@@
where $\Lambda^{\prime}_{k}=\max(0,-\lambda_{\min}(H_{k}))$, $c_{1}^{\prime},c_{2}^{\prime},\gamma^{\prime}$ are some positive parameter values and $t_{k}$ is the step-size chosen by Armijo’s step size rule, and show that this algorithm achieves $\left\|g_{k}\right\|\leq\varepsilon$ after at most $O(\varepsilon^{-2})$ iterations and it has a super-linear rate of convergence in a neighborhood of a local optimal solution under appropriate conditions.
@@
Let $\mathcal{D}$ denote the set of Gaussian matrices of size $s\times n$ whose entries are independently sampled from $\mathrm{N}(0,1/s)$. With a Gaussian random matrix $P_{k}$ from $\mathcal{D}$, the regularized sketched Hessian is defined by:
(6) $M_{k}:=P_{k}H_{k}P_{k}^{\mathsf{T}}+\eta_{k}I_{s}\in\mathbb{R}^{s\times s},$
@@
(7) $d_{k}:=-P_{k}^{\mathsf{T}}M_{k}^{-1}P_{k}g_{k}.$
The costly part of Newton-based methods, the inverse computation of a (approximate) Hessian matrix, is done in the subspace of size $s$. We note that $d_{k}$ defined by (7) is a descent direction for $f$ at $x_{k}$, i.e., $g_{k}^{\top}d_{k}<0$ if $g_{k}\neq 0$, since it turns out that $M_{k}$ is positive definite from the definition of $\Lambda_{k}$, and therefore $x^{\top}P_{k}^{\mathsf{T}}M_{k}^{-1}P_{k}x>0$ holds for $\forall x$ due to $P_{k}x\neq 0$ with high probability.
@@
(8) $f(x_{k})-f(x_{k}+\beta^{l_{k}}d_{k})\geq-\alpha\beta^{l_{k}}g_{k}^{\mathsf{T}}d_{k}.$
@@ 4 Global convergence properties
In Section 4.1, we discuss the global convergence of the RS-RNM under Section 4. We further prove the global iteration complexity of the algorithm in Section 4.2 by considering further assumptions.
@@
By Eq. 8, we have that for any $k\in\mathbb{N}$, $f(x_{k+1})\leq f(x_{k})$, implying all $x_{k}\in\Omega$. Since $\Omega$ is a bounded set and $f$ is continuously differentiable, there exists $U_{g}>0$ such that
(9) $\left\|g_{k}\right\|\leq U_{g},\ \forall k\geq 0.$
@@
(11) $\|H_{k}\|\leq L.$
Notice that by Eq. 10, $\nabla f$ is $L$-Lipschitz continuous. We also define $f^{*}=\inf_{x\in\Omega}f(x)$.
@@ 4.1 Global convergence
Suppose that $\left\|d_{k}\right\|\neq 0$. Then, $d_{k}$ defined by Eq. 7 satisfies
$\left\|d_{k}\right\|\leq\bar{\mathcal{C}}\frac{n}{s}\frac{\left\|g_{k}\right\|^{1-\gamma}}{c_{2}},$
@@
Proof. By Lemma 2.2 we have $\left\|P_{k}^{\mathsf{T}}P_{k}\right\|\leq\bar{\mathcal{C}}\frac{n}{s}$, holds with probability at least $1-2e^{-s}$. Then, it follows from Eq. 7 that
@@
We next show that, when $\left\|g_{k}\right\|$ is at least $\varepsilon$ away from 0, $\left\|d_{k}\right\|$ is bounded above by some constant.
@@
Suppose that Section 4 holds. Suppose also that there exists $\varepsilon>0$ such that $\left\|g_{k}\right\|\geq\varepsilon$. Then, with probability at least $1-2e^{-s}$, $d_{k}$ defined by Eq. 7 satisfies
(12) $\left\|d_{k}\right\|\leq r(\varepsilon),$
@@
Proof. If $\gamma\leq 1$, it follows from Lemma 4.1 and Eq. 9 that
@@
(13) $\left\|\nabla^{2}f(x)\right\|\leq U_{H}(\varepsilon),\ \forall x\in\Omega+B(0,r(\varepsilon)).$
@@
Suppose that Section 4 holds. Suppose also that there exists $\varepsilon>0$ such that $\left\|g_{k}\right\|\geq\varepsilon$. Then, with probability at least $1-2e^{-s}$, a step size $t_{k}^{\prime}>0$ such that
@@
$f(x_{k})-f(x_{k}+t_{k}^{\prime}d_{k})+\alpha t_{k}^{\prime}g_{k}^{\mathsf{T}}d_{k}$
(14) $=$ $(1-\alpha)t_{k}^{\prime}g_{k}^{\mathsf{T}}P_{k}^{\mathsf{T}}M_{k}^{-1}P_{k}g_{k}-\frac{1}{2}{t_{k}^{\prime}}^{2}g_{k}^{\mathsf{T}}P_{k}^{\mathsf{T}}M_{k}^{-1}P_{k}\nabla^{2}f(x_{k}+\tau_{k}^{\prime}t_{k}^{\prime}d_{k})P_{k}^{\mathsf{T}}M_{k}^{-1}P_{k}g_{k}$
(by (7))
[…]
(by (13))
@@
By Lemma 2.2, we have that, with probability at least $1-2e^{-s}$, $\left\|P_{k}P_{k}^{\mathsf{T}}\right\|\leq\frac{\bar{\mathcal{C}}n}{s}$. In addition, we have $\left\|H_{k}\right\|\leq U_{H}(\varepsilon)$ from Eq. 13, which gives us $\left\|P_{k}H_{k}P_{k}^{\mathsf{T}}\right\|\leq\frac{\bar{\mathcal{C}}n}{s}U_{H}(\varepsilon)$. For these reasons, we obtain evaluation of the values of $\lambda_{\min}(M_{k}^{-1})$ and $\lambda_{\max}(M_{k}^{-1}P_{k}P_{k}^{\mathsf{T}}M_{k}^{-1})$:
@@
(15) $\geq\frac{1}{\frac{\bar{\mathcal{C}}n}{s}U_{H}(\varepsilon)+c_{1}\frac{\bar{\mathcal{C}}n}{s}U_{H}(\varepsilon)+c_{2}\left\|g_{k}\right\|^{\gamma}},$
@@
(by (9) and $\left\|g_{k}\right\|\geq\varepsilon$)
@@
Suppose that Section 4 holds. Suppose also that there exists $\varepsilon>0$ such that $\left\|g_{k}\right\|\geq\varepsilon$. Then, with probability at least $1-2e^{-s}$, the step size $t_{k}$ chosen in Line 6 of RS-RNM satisfies
(16) $t_{k}\geq t_{\min}(\varepsilon),$
@@
we know that $t_{k}=1$ satisfies Armijo’s rule Eq. 8 from Lemma 4.5. If not, there exists $l_{k}\in\{0,1,2,\ldots\}$ such that
@@
and by Lemma 4.5, we have that the step size $\beta^{l_{k}+1}$ satisfies Armijo’s rule Eq. 8. Then, from the definition of $\beta^{l_{k}}$ in Line 6 of RS-RNM, we have
@@
Using Corollary 4.7, we can show the global convergence of RS-RNM under Section 4.
@@
Suppose that Section 4 holds. Let $\delta\in(0,1)$ and define $\delta_{s}:=2\left(\exp(-\frac{{\mathcal{C}}_{0}}{4}s)+\exp(-s)\right)$ and
$m=\left\lfloor\frac{f(x_{0})-f^{*}}{(1-\delta)(1-\delta_{s})p(\varepsilon)\varepsilon^{2}}\right\rfloor+1,\quad\text{where}\quad p(\varepsilon)=\frac{\alpha t_{\min}(\varepsilon)}{2\bar{\mathcal{C}}(1+c_{1})\frac{n}{s}U_{H}(\varepsilon)+2c_{2}U_{g}^{\gamma}}.$
Then, with probability at least $1-\exp\left(-\frac{\delta^{2}}{2}(1-\delta_{s})m\right)$ there exists $k\in\{0,1,\dots,m-1\}$ such that $\left\|g_{k}\right\|<\varepsilon$.
@@
Proof. We first notice that, by Lemma 2.1, applied with $\varepsilon=1/2$, and Lemma 2.2, we have, using Eq. 4, that $\left\|P_{k}g_{k}\right\|^{2}\geq\frac{1}{2}\left\|g_{k}\right\|^{2}$ and $\|P_{k}P_{k}^{\top}\|\leq\bar{\mathcal{C}}\frac{n}{s}$ holds for all $k\in\{0,1,\ldots,m-1\}$ with the given probability.
Suppose, for the sake of contradiction, that $\left\|g_{k}\right\|\geq\varepsilon$ for all $k\in\{0,1,\dots,m-1\}$. From Armijo’s rule Eq. 8, we can estimate how much the function value decreases in one iteration. We have that with probability $1-2\left(\exp(-\frac{{\mathcal{C}}_{0}}{4}s)+\exp(-s)\right)$:
@@
Let us denote by $\mathcal{A}_{k}$ the event, only depending of $P_{k}$, where the above inequality holds. Conditionally to the complement of $\mathcal{A}_{k}$ we have only that $f(x_{k})-f(x_{k+1})\geq 0$. Let us denote by $T_{k}\in\{0,1\}$ the random variable equal to $1$ if and only if $\mathcal{A}_{k}$ holds. Notice that the random variables $\{T_{k}\}$ are mutually independent because $T_{k}$ depends only on $P_{k}$. By the above remark we have that for all $k$: $f(x_{k})-f(x_{k+1})\geq p(\varepsilon)\varepsilon^{2}T_{k}$. Hence by adding up all these inequalities from $k=0$ to k=m1k=m-1, we get
(17) $f(x_{0})-f(x_{m})\geq p(\varepsilon)\varepsilon^{2}\sum\limits_{k=0}^{m-1}T_{k}.$
Since, for all $k$, $\mathbb{E}[T_{k}]\geq 1-2\left(\exp(-\frac{{\mathcal{C}}_{0}}{4}s)+\exp(-s)\right):=1-\delta_{s}$, we have by a Chernoff bound (see [41]), that for all $\delta\in(0,1)$,
(18) $\mathbb{P}\left(\sum\limits_{k=0}^{m-1}T_{k}\geq(1-\delta)(1-\delta_{s})m\right)\geq 1-\exp\left(-\frac{\delta^{2}}{2}(1-\delta_{s})m\right).$
Notice that by definition of $m$, we have that
$m>\frac{f(x_{0})-f^{*}}{(1-\delta)(1-\delta_{s})p(\varepsilon)\varepsilon^{2}}.$
Hence
(19) $(1-\delta)(1-\delta_{s})p(\varepsilon)\varepsilon^{2}m>f(x_{0})-f^{*}.$
Thus, with probability at least $1-\exp\left(-\frac{\delta^{2}}{2}(1-\delta_{s})m\right)$
@@
$\geq$ $(1-\delta)(1-\delta_{s})mp(\varepsilon)\varepsilon^{2}$
$>$ $f(x_{0})-f^{*},$
where the second inequality holds by (17) together with (18) and the strict inequality holds by Eq. 19. This is a contradiction, hence there exists $k\in\{0,1,\dots,m-1\}$ such that $\left\|g_{k}\right\|<\varepsilon$.